OpenStudy (anonymous):

use the distributive property to simplify thee expression. -7(8x-3)+x A.x+10 B.-15x+4 C.-9x-4 D.-55x+21 HELPPLZ

OpenStudy (anonymous):

a(b+c) = ab+ac ... this is the distributive property.

OpenStudy (anonymous):

use it on the -7(8x-3) part and then combine any like terms.

OpenStudy (akashdeepdeb):

This is how the distributive property goes. x times (a+b) means x times a + x times b OR x(a+b) = xa+xb So -7(8x-3)+x -7*8x + 7*3 + x Now solve it and get your answer. Getting me? :)

OpenStudy (anonymous):

who wants a metal and a fan ?

OpenStudy (anonymous):

is it D OR c?

OpenStudy (anonymous):

C